# Suiren-1.0 Technical Report: A Family of Molecular Foundation Models

### 23 Mar 2026 v1

@@ Appendix C Evaluation of MoleHB Scaffold split
  • Energetic properties (Table 21): This category exhibits the most pronounced performance disparity. Baseline methods suffer severe degradation (e.g., MoleBERT’s MAE for enthalpy of combustion exceeds 4400), whereas Suiren-ConfAvg maintains substantially lower errors (545.71), representing an 82.8% relative improvement. We hypothesize that the pre-training objective of Suiren-Base, which incorporates physics-informed energy constraints, enables more transferable representations for thermodynamic quantities that are sensitive to subtle electronic and conformational features.
  • Critical and saturation properties (Table 15): Suiren-ConfAvg dominates in critical temperature (21.75 vs. 63.35, best baseline) and critical volume (59.15 vs. 290.13), with relative improvements of 65.7% and 79.6%, respectively. However, for critical density and critical pressure, MoleBERT achieves marginally better results (0.00671 vs. 0.00690 and 1.737 vs. 1.932). This suggests that certain intensive properties with strong linear correlations to molecular size may be adequately captured by simpler architectures, whereas extensive or composite properties benefit from Suiren’s enhanced representation capacity.
  • Thermal and fluctuation properties (Tables 1917): The density of liquid shows minimal performance variation across splits (Suiren-ConfAvg: 0.01023), consistent with its strong dependence on atomic composition rather than scaffold topology. In contrast, heat capacities and thermal expansion coefficients exhibit substantial scaffold sensitivity, where Suiren-ConfAvg achieves 59.7%–86.8% relative improvements.
  • Structural properties (Table 20): While Suiren-ConfAvg excels in liquid volume (70.9% improvement) and radius of gyration (60.3% improvement), it is slightly outperformed by Uni-Mol v1 on refractive index and dipole moment.
  • Safety, solution, and transport properties (Tables 1622): Suiren-ConfAvg consistently attains the best or near-best performance, with particularly notable gains in flash point (36.7% improvement) and vapor pressure (56.2% improvement).

### 13 Apr 2026 v2

@@ Appendix C Evaluation of MoleHB Scaffold split
  • Energetic properties (Table 21): This category exhibits the most pronounced performance disparity. Baseline methods suffer severe degradation (e.g., MoleBERT’s MAE for enthalpy of combustion is 4411.49), whereas Suiren-ConfAvg maintains substantially lower errors (338.81), representing an 89.32% relative improvement. We hypothesize that the pre-training objective of Suiren-Base, which incorporates physics-informed energy constraints, enables more transferable representations for thermodynamic quantities that are sensitive to subtle electronic and conformational features.
  • Critical and saturation properties (Table 15): Suiren-ConfAvg dominates in critical temperature (19.59 vs. 63.35, best baseline) and critical volume (55.80 vs. 290.13), with relative improvements of 69.07% and 80.77%, respectively. However, for critical density, MoleBERT achieves marginally better results (0.00671 vs. 0.00888). For critical pressure, Suiren achieves 1.66 vs. MoleBERT’s 1.74, with a 4.68% improvement. This suggests that certain intensive properties with strong linear correlations to molecular size may be adequately captured by simpler architectures, whereas extensive or composite properties benefit from Suiren’s enhanced representation capacity.
  • Thermal and fluctuation properties (Tables 1917): The density of liquid shows moderate performance variation across splits (Suiren-ConfAvg: 0.00851), consistent with its dependence on atomic composition, with a 29.67% improvement over the best baseline. In contrast, heat capacities exhibit substantial scaffold sensitivity, with Suiren-ConfAvg achieving 41.50%–88.63% improvements for liquid and solid heat capacities respectively, while showing 72.33% improvement for gas heat capacity.
  • Structural properties (Table 20): While Suiren-ConfAvg excels in radius of gyration (57.65% improvement), it is slightly outperformed by Uni-Mol v1 on refractive index (0.00859 vs. 0.00673) and dipole moment (0.49490 vs. 0.57567 for Suiren).
  • Safety, solution, and transport properties (Tables 1622): Suiren-ConfAvg consistently attains the best or near-best performance, with particularly notable gains in flash point (28.99% improvement) and vapor pressure (47.30% improvement).
@@
Appendix D Ablation Studies of Suiren-ConfAvg
To systematically assess the contributions of Suiren-ConfAvg and DGNN to downstream task performance, we designed a comprehensive ablation study with four distinct model variants: (1) Baselines: the complete architecture with DGNN+Suiren-ConfAvg serving as our baseline; (2) Suiren-ConfAvg only ( ): the DGNN component removed, with Suiren-ConfAvg acting as the sole backbone—prediction heads are attached to frozen Suiren-ConfAvg embeddings, allowing only the heads to be trained; (3) Suiren-ConfAvg only ( ): built upon variant (2) but with Suiren-ConfAvg weights unfrozen and co-optimized end-to-end; (4) w/o. Suiren-ConfAvg: Suiren-ConfAvg pre-trained weights entirely discarded, replaced by a deep task-specific GNN trained from scratch. To ensure the generalizability of our experimental findings, we selected properties spanning diverse domains to highlight the representational differences across distinct tasks.
Table 23: Results of ablation studies on MoleHB random split.
Property Baselines Suiren-ConfAvg Suiren-ConfAvg w/o.
only ( ) only ( ) Suiren-ConfAvg
Critical Temperature 6.6725 8.6190 7.3257 7.9764
Flash Point 8.2072 9.5567 8.8750 9.2798
Melting Point 11.0789 12.7397 12.4325 12.7382
Surface Tension 0.8205 0.9020 0.8568 0.9123
From the results presented in Table 23, we draw the following conclusions: (1) Necessity of task-specific adaptation: The removal of the task-specific GNN from the DGNN architecture precipitates significant performance degradation, with this effect being particularly pronounced when Suiren-ConfAvg weights remain frozen. This finding underscores that effective transfer of Suiren-ConfAvg representations to downstream tasks necessitates sufficient learnable capacity and modular adaptation, while simultaneously corroborating the existence of substantial distributional divergence across macroscopic property prediction tasks. (2) Efficacy of pre-trained representations: The exclusion of Suiren-ConfAvg pre-trained weights uniformly impairs performance across all evaluated tasks, yielding inferior results compared to the Suiren-ConfAvg-with-prediction-head configuration. This establishes the beneficial role of Suiren-ConfAvg representations in enhancing downstream predictive capabilities.
Interestingly, the task-specific GNN trained from scratch achieves competitive performance relative to prior molecular foundation models. This phenomenon may be attributed to two factors: (1) the inherent limitations of conventional molecular representations in capturing the broad spectrum of macroscopic physicochemical properties; (2) architectural innovations in our 2D GNN design, which integrates fully-connected GAT layers with adjacency-matrix-based GAT layers to enhance expressive capacity. These ablation studies provide compelling empirical evidence for the indispensability of synergistically combining Suiren-ConfAvg pre-trained representations with the adaptive capacity of the DGNN architecture.